OpenStudy (anonymous):

Solve for x: −2|x − 3| = −12

OpenStudy (anonymous):

|x-3|=6

OpenStudy (anonymous):

x=-3 x=9

OpenStudy (johnweldon1993):

You should first get the absolute value by itself....so divide both sides of the equation by -2 |x - 3| = 6 Now split this into 2 equations x - 3 = 6 and x - 3 = -6 And solve both equations for 'x'

hartnn (hartnn):

you can start by dividing both sides by -2 what u get ?